OpenStudy (anonymous):

The pressure (P) exerted on an object in the ocean at a depth (d) of 100 meters is 11 atmospheres. Supposing that the object is descending at a rate of 2 m/s, find the rate of change of pressure at a depth of 100 meters. (P = kd + 1, where k is a constant.)

OpenStudy (anonymous):

\[P(t)=kd(t)+1\] \[P'(t)=kd'(t)\] But \[d'(t)=2\] So \[P'(t)=2k\] for any depth.

OpenStudy (anonymous):

from P=kd+1 11atm =k(100)+1, k=10/100=0.1= constant DP(t)=kDd(t), Dd(t)=2m/s, DP(t)=kDd(t)=(0.1)(2) = 0.2 atmospheric pressure per sec ans